Joan is 5 years older than Ellen, and 3 years ago the sum of their ages was 17 years. In how many years will Joan be 21 years old?
Answer by ankor@dixie-net.com(22740)

Joan is 5 years older than Ellen,
j = e + 5
j - e = 5
and 3 years ago the sum of their ages was 17 years.
(j-3) + (e-3) = 17
j + e - 6 = 17
j + e = 17 + 6
j + e = 23
use elimination on these two equations to find j
j - e = 5
j + e = 23
-------------Addition eliminates e, find j
2j = 28
j = 14 yrs
In how many years will Joan be 21 years old?
Obviously she will be 21 in 7 yrs